500 milliliters of 20% acid solution needs to be combined with what percent solution to get 800 milliliters of 38 solution?
Answer by stanbon(75887)

500 milliliters of 20% acid solution needs to be combined with what percent solution to get 800 milliliters of 38% solution?
-----------
20% solution DATA:
Amt = 500 ml ; amt of acid is 0.20*500 = 100 ml
------------
x% solution DATA:
Amt. = 300 ml ; amt of acid = (x/100)*300 = 3x ml
38% solution DATA:
Amt = 800 ml ; Amt of acid = 0.38*800 = 304 ml
----------------
EQUATION:
acid + acid = acid
100 + 3x = 304
3x = 204
x = 68% (the 500 ml has to be combined with 300 ml of 68% solution)
============
Cheers,
Stan H.